OpenStudy (ksaimouli):

zanchary the circus clown is shot at 65 degree with horizontal.how fast should zanchary be moving if he is to land in the middle of the net 22m away?assume the net is placed at the same level

OpenStudy (ksaimouli):

can any one help

OpenStudy (anonymous):

 \[v_0(y)=v_0\sin \theta\]\[v_0(x)=v_0\cos \theta\] \[s_y(t)=v_0\sin \theta-\frac{1}{2}g t^2\]\[s_x(t)=v_0\cos \theta\]

OpenStudy (ksaimouli):

ok

OpenStudy (anonymous):

*\[s_x(t)=v_0\cos \theta t\] and add a t term for s_y

OpenStudy (ksaimouli):